# Boosting Memory Efficiency in Transfer Learning for High-Resolution Medical Image Classification

### v2

@@ IV-C Comparisons with State-of-the-art
We compare FPT+ against full fine-tuning, linear probing, and SOTA PETL approaches, including our previous work FPT. In full fine-tuning, all parameters are learnable when training on downstream tasks. Although full fine-tuning often represents the performance’s upper bound for fine-tuning, the relative scarcity of training samples in medical image datasets may lead to overfitting and consequent performance degradation. Linear probing emerges as an efficient strategy to adapt pre-trained models to downstream tasks by solely training a new task-specific head on top of the pre-trained model. Additionally, we compare FPT+ with popular SOTA PETL methods applicable to vision tasks.
@@
Although these aforementioned methods achieve reductions in the number of learnable parameters, their memory consumption remains high. For instance, Adapter has the lowest memory consumption but still requires 19,360 MB to fine-tune the LPM. In contrast, our proposed method, FPT+, outperforms other PETL methods in both performance and efficiency, achieving the best performance among PETL methods for both pre-trained weights and the best performance-efficiency balance (PPE and PME) among all methods. Specifically, FPT+ exhibits remarkable parameter efficiency with only 1.03% of learnable parameters and significantly lower memory consumption (736 MB, 3.18% compared to full fine-tuning), without compromising performance. Our previous work, FPT, already shows significant improvements in parameter efficiency and performance. On top of FPT, FPT+ further enhances these metrics, achieving higher performance and reduced memory consumption (736 MB vs. 1,824 MB).
The comparison results emphasize the effectiveness of FPT+ in balancing performance, parameter efficiency, and memory consumption. Its design ensures efficient high-resolution medical image classification, making it a promising solution for advancing medical diagnostics. The generalizability and robustness of FPT+ are further highlighted by its effectiveness with both supervised and DINO pre-trained weights, making it adaptable to different pre-training paradigms and practical for a wide range of medical image classification tasks.
@@ IV-D Training Efficiency
As shown in Table II, FPT+ demonstrates remarkable training efficiency with computational cost of only 0.06 GB FLOPs, being significantly lower than full fine-tuning (87.84 GB) and PETL methods like LoRA (107.89 GB). FPT+ also excels in terms of throughput, processing 357.37 images per second, which is a significant improvement over other PETL methods like LoRA (20.62 images/second) and Prompt-tuning (22.19 images/second). Compared to FPT, FPT+ doubles the throughput, showing its enhanced efficiency. With these reductions in the computational cost, FPT+ achieves a total training time of 90.61 seconds, which is more than 12 times faster than full fine-tuning (1135.91 seconds). FPT+ is also energy-efficient, achieving an average power usage of 107.60 W and a peak power usage of 114.38 W, staying far below the average of 280 W and the peak usage of over 290 W for other PETL methods on GPUs with a 300 W power limit.
@@
IV-E Evaluation with A Larger Pre-trained Model
To further investigate the generalizability of FPT+ and highlight its significant advantage in memory saving, we evaluate its performance and efficiency by comparing it with SOTA methods using a ViT-L backbone with supervised pre-trained weights across different input resolutions. The superiority of FPT+ becomes even more evident when using a larger model. As shown in Table III, most PETL methods encounter out-of-memory issues on a 48GB GPU when the input resolution exceeds 384 $\times$ 384, limiting their capability to improve performance through higher resolutions. In contrast, FPT, FPT+, and linear probing, with better memory efficiency, can run at higher resolutions and exhibit greater performance potential.
@@ IV-E Evaluation with A Larger Pre-trained Model
IV-F Contribution of Each Layer
In this section, we analyze the behavior of our previous work, FPT, in terms of fine-grained prompts and show how FPT+ empirically simplifies FPT’s architecture. As described in FPT, similar to FPT+, the pre-trained knowledge is conveyed to the side network via fine-grained prompts and fusion modules. The fine-grained prompts are concatenated with the intermediate sequence of the side network to join the forward propagation. The side network integrates the pre-trained knowledge through self-attention layers, where other tokens obtain information from the fine-grained prompts by paying attention to these prompts. However, unlike FPT+, which has a side network of $L_{S}$ layers, in FPT, each layer in the LPM has a corresponding learnable layer and a corresponding fusion module, resulting in a side netowrk of $L_{M}$ layers, which induces non-trivial computational overhead.
@@ IV-G Ablation Study
IV-G1 Impact of Components
To assess the contributions of individual components within our proposed FPT+, we conduct an ablation study by adding components sequentially and measuring their impact on AUC and GPU memory usage. As shown in Fig. 7, each additional component incrementally enhances the model’s performance, with the average AUC improving from 74.74 to 87.12. The most substantial performance gain is attributed to the introduction of pre-trained knowledge from the LPM. However, simply using MLP for transferring pre-trained knowledge significantly increases the memory consumption from 8,214 MB to 35,682 MB. Replacing MLP with fusion modules notably enhances memory efficiency by reducing the memory consumption from 35,682 MB to 13,328 MB, followed by consistent improvements with the incorporation of asymmetric inputs, important tokens selection, and features preloading down to 736 MB. By combining these proposed components, FPT+ achieves the optimal balance of high performance and low memory consumption.
@@ IV-G1 Impact of Components
IV-G2 Input Resolutions
To reduce memory consumption, we propose asymmetric inputs for the LPM and the side network, where the input for the side network is a down-scaled version of that for the LPM. In this section, we investigate the impact of the input resolution for the side network. We train FPT+ using different input resolutions for the side network while maintaining the input resolution for the LPM at $512\times 512$. As shown in Table IV, the results indicate that a resolution of $128\times 128$ yields optimal performance while maintaining memory efficiency. We conjecture that the poor performance observed with a higher down-scaled resolution could potentially be attributed to the lengthy input sequence in the side network, which distracts attention from the fine-grained prompts. Notably, even using an input resolution of $64\times 64$ for the side network achieves better performance compared to other PETL methods, demonstrating that the proposed fine-grained prompts and fusion modules can effectively convey fine-grained information for diagnostics in the side network.
TABLE IV: FPT+’s performance and efficiency with different input resolutions for the side network.
Input Resolution Avg. AUC $\uparrow$ Mem. $\downarrow$ PME $\uparrow$
@@ IV-G2 Input Resolutions
IV-G3 Number of Fine-grained Prompts
The number of fine-grained prompts can be significantly fewer than the total tokens from the input images presented in both the pre-trained model and the side network, thereby reducing the memory consumption. We here explore the influence of the number of fine-grained prompts. As depicted in Table V, with the increase in the number of prompts, the performance improves and the best performance is achieved at 16 prompts, which number is sufficient to summarize fine-grained information and convey it to the side network. We also notice that even with only 4 prompts, the performance is comparable to other PETL methods, demonstrating the effectiveness of FPT+’s fusion modules in obtaining pre-trained knowledge from the LPM.
TABLE V: FPT+’s performance and efficiency with different numbers of prompts for the side network.
@@ IV-G3 Number of Fine-grained Prompts
IV-G4 Reduction Factor
The reduction factor $k$ in FPT+ determines the hidden dimension of the side network. As shown in Table VI, we assess the impact of different reduction factors on FPT+’s performance and efficiency. A reduction factor of 8 achieves the highest average AUC (87.12) with minimal memory usage (736 MB) and only 1.03% of learnable parameters. Increasing the reduction factor to 16 slightly reduces memory usage but results in a drop in AUC (85.92), while lower factors (e.g., 2 and 4) increase resource consumption without significant gains in performance. We observe that a larger hidden dimension in the side network (lower reduction factor) does not improve performance as it increases the training complexity of the randomly initialized side network. A reduction factor of 8 achieves the optimal balance for high-resolution medical image classification tasks.
TABLE VI: FPT+’s performance and efficiency with different reduction factors.
Reduction Factor Avg. AUC $\uparrow$ Param. $\downarrow$ Mem. $\downarrow$ PPE $\downarrow$ PME $\downarrow$
@@ IV-G4 Reduction Factor
IV-G5 Ratio of Important Tokens
To mitigate the training cost, we exclusively integrate features derived from important tokens of the pre-trained model into the side network. In this section, we assess how different ratios of these important tokens affect the process. We provide details on the average performance and memory usage of FPT+ across various token ratios. We notice that the optimal ratio of important tokens varies across different datasets. For instance, although both DDR and Messidor-2 are fundus image datasets, they show peak performance at different important token ratios. Although a higher ratio of important tokens can provide more information for fusion modules, it carries a risk of misleading the fusion modules and potentially increasing the difficulty of learning. Most datasets prefer a token ratio lower than 0.5, except for Messidor-2 and CHNCXR which achieve the best performance at 0.9 and 0.7, respectively. We conjecture that the optimal token ratio depends on the scale and complexity of the dataset of interest.
In terms of memory efficiency, as the ratio of important tokens increases, the memory consumption also rises dramatically, resulting in significant drops in PME. Overall, a token ratio of 0.2 emerges as the optimal choice. At this ratio, FPT+ achieves the best average AUC with low memory consumption, ensuring an efficient balance. The PME metric also indicates that a 0.2 ratio provides the best balance between performance and memory efficiency.
@@ IV-G5 Ratio of Important Tokens
Figure 8: Impact of important tokens’ ratio on FPT+’s performance and efficiency. The star symbol indicates the best performance. Increasing the ratio does not affect the number of learnable parameters, resulting in a constant PPE (represented as a horizontal line).
IV-H Visualization

### v3

@@ IV-B2 Evaluation Metrics
ViT-B
Full FT 2,716 4,540 10,034 23,128 48,268 83.63 86.28 87.50 88.82 89.11
Linear probing 842 1,416 1,668 3,416 6,836 77.81 82.06 82.69 83.28 82.62
Prompt-tuning [27] 1,266 3,094 8,204 20,582 45,352 80.03 82.68 83.62 83.69 83.33
Attention-tuning [45] 1,584 3,492 8,300 20,792 45,396 79.98 82.51 82.94 83.68 82.35
Adapter [7] 1,280 3,018 7,962 19,360 43,516 80.94 81.61 82.80 84.14 84.69
BitFit [2] 1,250 2,964 8,072 20,382 45,392 80.80 84.87 85.34 85.67 86.96
LoRA [22] 1,294 3,132 8,472 20,970 46,554 81.93 84.67 85.48 85.94 86.96
AdaLoRA [51] 1,290 3,120 8,360 20,960 45,798 81.10 84.86 85.97 86.59 87.11
FPT [23] 1,430 1,522 1,682 1,824 2,166 81.22 84.92 85.63 86.40 86.19
FPT+ (Ours) 549 624 698 736 896 81.29 85.07 86.02 87.12 87.54
ViT-L
@@ IV-C Comparisons with State-of-the-art
We compare FPT+ against full fine-tuning, linear probing, and SOTA PETL approaches. In full fine-tuning, all parameters are learnable when training on downstream tasks. Although full fine-tuning often represents the performance’s upper bound for fine-tuning, the relative scarcity of training samples in medical image datasets may lead to overfitting and consequent performance degradation. Linear probing emerges as an efficient strategy to adapt pre-trained models to downstream tasks by solely training a new task-specific head on top of the pre-trained model. Additionally, we compare FPT+ with popular SOTA PETL methods applicable to vision tasks.
@@
Although these aforementioned methods achieve reductions in the number of learnable parameters, their memory consumption remains high. For instance, Adapter has the lowest memory consumption but still requires 19,360 MB to fine-tune the LPM. In contrast, our proposed method, FPT+, outperforms other PETL methods in both performance and efficiency, achieving the best performance among PETL methods and the best performance-efficiency balance (PPE and PME) among all methods across different pre-trained weights. Specifically, FPT+ exhibits remarkable parameter efficiency with only 1.03% of learnable parameters and significantly lower memory consumption (736 MB, 3.18% compared to full fine-tuning), without compromising performance. FPT already shows significant improvements in parameter efficiency and performance. On top of FPT, FPT+ further enhances these metrics, achieving higher performance and reduced memory consumption (736 MB vs. 1,824 MB). Notably, the memory consumption reported for FPT is lower than that reported in its original paper, due to an optimized preloading mechanism in FPT+ that stores only the important tokens from the frozen LPM. This optimization, applied to both FPT+ and FPT, reduces memory usage without affecting the AUC performance.
In addition to using ImageNet-based pre-trained weights, we evaluate FPT+ with BioMedCLIP [52], a widely used vision-language model pre-trained on a diverse set of medical imaging modalities. In this evaluation, the vision encoder of BioMedCLIP is adopted as the LPM. As shown in Table I, FPT+ with BioMedCLIP consistently outperforms other PETL methods across five datasets, achieving the highest average AUC. Notably, it surpasses the full fine-tuning baseline in terms of the average AUC, further highlighting its ability to effectively leverage medical-specific pre-trained knowledge.
The comparison results emphasize the effectiveness of FPT+ in balancing performance, parameter efficiency, and memory consumption. Its design ensures efficient high-resolution medical image classification, making it a promising solution for advancing medical diagnostics. The generalizability and robustness of FPT+ are further highlighted by its effectiveness with different pre-trained weights, making it adaptable to different pre-training paradigms and practical for a wide range of medical image classification tasks.
@@ IV-D Training Efficiency
As shown in Table II, FPT+ demonstrates remarkable training efficiency with computational cost of only 0.06 GB FLOPs, being significantly lower than full fine-tuning (87.84 GB) and PETL methods like LoRA (107.89 GB). FPT+ also excels in terms of throughput, processing 357.37 images per second, which is a significant improvement over other PETL methods like LoRA (20.62 images/second) and Prompt-tuning (22.19 images/second). Compared to FPT, FPT+ doubles the throughput, showing its enhanced efficiency. With these reductions in the computational cost, FPT+ achieves a total training time of 90.61 seconds, which is more than 12 times faster than full fine-tuning (1,135.91 seconds). FPT+ is also energy-efficient, achieving an average power usage of 107.60 W and a peak power usage of 114.38 W, staying far below the average of 280 W and the peak usage of over 290 W for other PETL methods on GPUs with a 300 W power limit.
@@
IV-E Evaluation with Different Resolutions
To further validate the resolution scalability and memory efficiency of FPT+, we compare its performance against other PETL baselines using the ViT-B backbone with ImageNet-based supervised LPM across five input resolutions: 128 × 128, 256 × 256, 384 × 384, 512 × 512, and 640 × 640.
As shown in the top panel of Table III, FPT+ consistently achieves strong AUC performance across all resolutions. Notably, it achieves the best average AUC across all resolutions except for the lowest resolution (128 × 128), where it still demonstrates competitive performance compared to other PETL methods. Importantly, FPT+ demonstrates exceptional memory efficiency. At a resolution of 640 × 640, FPT+ consumes only 896 MB, which is significantly lower than all competing PETL methods (e.g., 46,554 MB for LoRA). This highlights the scalability of FPT+ in handling high-resolution inputs with minimal computational overhead.
These experimental results also confirm that FPT+ not only scales effectively across a wide range of resolutions but also achieves an optimal balance between performance and efficiency, making it well-suited for deployment in various clinical settings.
TABLE IV: Performance of FPT+ using ViT-B with a patch size of 8 and the ImageNet-based supervised LPM at a resolution of 256 × 256.
Full FT 100 23,078 87.76 $\pm$ 0.79 84.95 $\pm$ 0.99 96.52 $\pm$ 0.06 81.87 $\pm$ 0.69 93.18 $\pm$ 0.45 67.26 $\pm$ 1.76 99.85 $\pm$ 0.03 96.27 $\pm$ 0.36 88.46 65.47 65.47
Linear probing 0.01 3,366 81.55 $\pm$ 0.48 82.34 $\pm$ 0.36 93.65 $\pm$ 0.04 80.94 $\pm$ 0.62 81.94 $\pm$ 0.34 59.89 $\pm$ 1.54 99.23 $\pm$ 0.11 90.11 $\pm$ 0.42 83.71 83.71 78.90
Prompt-tuning [27] 0.17 20,530 81.72 $\pm$ 0.56 82.26 $\pm$ 0.31 93.84 $\pm$ 0.10 81.31 $\pm$ 0.70 83.03 $\pm$ 0.32 59.83 $\pm$ 1.01 99.25 $\pm$ 0.09 92.78 $\pm$ 0.30 84.25 84.19 63.91
Attention-tuning [45] 32.97 20,740 83.28 $\pm$ 1.46 82.92 $\pm$ 1.30 94.28 $\pm$ 0.65 78.39 $\pm$ 2.25 82.94 $\pm$ 3.06 60.17 $\pm$ 5.15 99.79 $\pm$ 0.06 92.21 $\pm$ 0.67 84.25 74.44 63.77
Adapter [7] 2.04 19,596 88.33 $\pm$ 0.82 81.69 $\pm$ 1.68 93.69 $\pm$ 0.28 78.39 $\pm$ 3.09 81.81 $\pm$ 3.24 61.19 $\pm$ 0.44 99.84 $\pm$ 0.03 89.96 $\pm$ 1.85 84.36 83.62 64.59
BitFit [2] 0.12 20,330 85.02 $\pm$ 0.24 84.69 $\pm$ 0.32 93.92 $\pm$ 0.10 82.20 $\pm$ 1.60 83.44 $\pm$ 0.26 65.17 $\pm$ 1.54 99.68 $\pm$ 0.07 92.57 $\pm$ 0.67 85.84 85.80 65.24
LoRA [22] 0.69 20,936 86.38 $\pm$ 0.47 85.03 $\pm$ 0.83 94.44 $\pm$ 0.22 79.96 $\pm$ 3.40 80.34 $\pm$ 2.59 59.78 $\pm$ 1.08 99.84 $\pm$ 0.04 93.77 $\pm$ 0.76 84.94 84.69 64.17
AdaLoRA [51] 0.52 20,926 86.24 $\pm$ 0.35 84.78 $\pm$ 0.55 94.41 $\pm$ 0.22 82.95 $\pm$ 1.14 87.13 $\pm$ 0.50 64.11 $\pm$ 0.88 99.50 $\pm$ 0.12 93.50 $\pm$ 0.52 86.58 86.39 65.42
FPT [23] 1.83 9,144 84.07 $\pm$ 0.54 84.27 $\pm$ 0.89 94.33 $\pm$ 0.64 82.16 $\pm$ 1.82 89.35 $\pm$ 0.74 63.85 $\pm$ 0.52 98.77 $\pm$ 0.41 89.35 $\pm$ 2.85 85.77 85.10 74.20
FPT+ (Ours) 0.99 1,494 85.48 $\pm$ 0.74 86.95 $\pm$ 0.41 94.50 $\pm$ 0.30 83.93 $\pm$ 1.06 89.85 $\pm$ 1.25 67.46 $\pm$ 0.45 99.54 $\pm$ 0.20 90.28 $\pm$ 2.24 87.25 86.88 84.91
IV-F Evaluation with A Larger Pre-trained Model
To further assess the generalizability of FPT+ and highlight its significant advantage in memory saving when scaling model capacity, we evaluate its performance using the ViT-L backbone across various input resolutions. As shown in Table III, most PETL methods encounter out-of-memory issues on a 48GB GPU when the input resolution exceeds 384 $\times$ 384, limiting their capability to improve performance through higher resolutions. In contrast, FPT, FPT+, and linear probing, with better memory efficiency, can run at higher resolutions and exhibit greater performance potential.
@@ IV-F Evaluation with A Larger Pre-trained Model
IV-G Evaluation with Smaller Patch Size
To further explore the adaptability of FPT+ to different ViT configurations, we evaluate its performance using ViT-B with a smaller patch size of 8 at an input resolution of 256 × 256. Although a smaller patch size can help the model capture fine-grained information, it substantially increases the number of input tokens, leading to significantly higher computational costs. As only FPT+, FPT, and linear probing are feasible at a resolution of 512 × 512 on a 48GB GPU under this configuration, We conduct experiments at a resolution of 256 × 256 for all methods to enable a fair comparison.
As shown in Table IV, FPT+ maintains remarkable efficiency under this memory-demanding setup. While most PETL baselines consume approximately 20,000 MB of memory, FPT+ requires only 1,494 MB, a footprint comparable to that of ViT-B with a patch size of 16 at a 512 × 512 resolution. Despite its significantly lower memory usage, FPT+ achieves the best AUC performance on five out of eight evaluated datasets, along with the highest average AUC and performance-efficiency balance among all methods. These results further confirm the robustness of FPT+ across various ViT configurations.
IV-H Contribution of Each Layer
In this section, we analyze the behavior of FPT, in terms of fine-grained prompts and show how FPT+ empirically simplifies FPT’s architecture. As described in FPT, similar to FPT+, the pre-trained knowledge is conveyed to the side network via fine-grained prompts and fusion modules. The fine-grained prompts are concatenated with the intermediate sequence of the side network to join the forward propagation. The side network integrates the pre-trained knowledge through self-attention layers, where other tokens obtain information from the fine-grained prompts by paying attention to these prompts. However, unlike FPT+, which has a side network of $L_{S}$ layers, in FPT, each layer in the LPM has a corresponding learnable layer and a corresponding fusion module, resulting in a side netowrk of $L_{M}$ layers, which induces non-trivial computational overhead.
@@ IV-I Ablation Study
IV-I1 Impact of Components
To assess the contributions of individual components within our proposed FPT+, we conduct an ablation study by adding components sequentially and measuring their impact on AUC and GPU memory usage. As shown in Fig. 7, each additional component incrementally enhances the model’s performance, with the average AUC improving from 74.74 to 87.12. The most substantial performance gain is attributed to the introduction of pre-trained knowledge from the LPM. However, simply using MLP for transferring pre-trained knowledge significantly increases the memory consumption from 8,214 MB to 35,682 MB. Replacing MLP with fusion modules notably enhances memory efficiency by reducing the memory consumption from 35,682 MB to 13,328 MB, followed by consistent improvements with the incorporation of asymmetric inputs, important tokens selection, and features preloading down to 736 MB. Notably, important tokens selection mechanism not only lowers memory consumption but also improves the AUC performance by reducing the number of non-informative tokens, allowing the fusion module to focus on diagnostically relevant features. By combining these components, FPT+ achieves the optimal balance of high performance and low memory consumption.
@@ IV-I1 Impact of Components
TABLE V: FPT+’s performance and efficiency with different input resolutions for the side network.
Side Input Resolution Avg. AUC $\uparrow$ Mem. $\downarrow$ PME $\uparrow$
@@
IV-I2 Side Input Resolutions
To reduce memory consumption, we propose asymmetric inputs for the LPM and the side network, where the input for the side network is a down-scaled version of that for the LPM. In this section, we investigate the impact of the input resolution for the side network. We train FPT+ using different input resolutions for the side network while maintaining the input resolution for the LPM at $512\times 512$. As shown in Table V, the results indicate that a resolution of $128\times 128$ yields optimal performance while maintaining memory efficiency. We conjecture that the poor performance observed with a higher down-scaled resolution could potentially be attributed to the lengthy input sequence in the side network, which distracts attention from the fine-grained prompts. Notably, even using an input resolution of $64\times 64$ for the side network achieves better performance compared to other PETL methods, demonstrating that the proposed fine-grained prompts and fusion modules can effectively convey fine-grained information to the side network.
IV-I3 Number of Fine-grained Prompts
The number of fine-grained prompts can be significantly fewer than the total tokens from the input images presented in both the pre-trained model and the side network, thereby reducing the memory consumption. We here explore the influence of the number of fine-grained prompts. As depicted in Table VI, with the increase in the number of prompts, the performance improves and the best performance is achieved at 16 prompts, which number is sufficient to summarize fine-grained information and convey it to the side network. We also notice that even with only 4 prompts, the performance is comparable to other PETL methods, demonstrating the effectiveness of FPT+’s fusion modules in obtaining pre-trained knowledge from the LPM.
TABLE VI: FPT+’s performance and efficiency with different numbers of prompts for the side network.
@@ IV-I3 Number of Fine-grained Prompts
TABLE VII: FPT+’s performance and efficiency with different reduction factors.
Reduction Factor Avg. AUC $\uparrow$ Param. $\downarrow$ Mem. $\downarrow$ PPE $\uparrow$ PME $\uparrow$
@@
TABLE VIII: FPT+’s performance with different token selection strategies.
Selection strategy Avg. AUC $\uparrow$
Random tokens selection 85.79
Important tokens selection 87.12
IV-I4 Reduction Factor
The reduction factor $k$ in FPT+ determines the hidden dimension of the side network. As shown in Table VII, we assess the impact of different reduction factors on FPT+’s performance and efficiency. A reduction factor of 8 achieves the highest average AUC (87.12) with minimal memory usage (736 MB) and only 1.03% of learnable parameters. Increasing the reduction factor to 16 slightly reduces memory usage but results in a drop in AUC (85.92), while lower factors (e.g., 2 and 4) increase resource consumption without significant gains in performance. We observe that a larger hidden dimension in the side network (lower reduction factor) does not improve performance as it increases the training complexity of the randomly initialized side network. A reduction factor of 8 achieves the optimal balance for high-resolution medical image classification tasks.
IV-I5 Token Selection Strategy
To validate the effectiveness of the proposed important token selection strategy, we conduct an ablation study comparing it with a random token selection baseline. In the random strategy, tokens are selected uniformly at random rather than guided by attention-based importance scores. As presented in Table VIII, important tokens selection consistently outperforms the random counterpart, achieving an average improvement of 1.33% in AUC across all evaluated datasets. These results confirm that selecting diagnostically relevant tokens based on attention scores enables more efficient and effective transfer of pre-trained knowledge.
@@ IV-I5 Token Selection Strategy
Figure 8: Impact of important tokens’ ratio on FPT+’s performance and efficiency with a 512 × 512 input resolution. The star symbol indicates the best one.
Figure 9: Impact of important tokens’ ratio on FPT+’s performance and efficiency with a 256 × 256 input resolution. The star symbol indicates the best one.
IV-I6 Ratio of Important Tokens
To mitigate the training cost, we exclusively integrate features derived from important tokens of the pre-trained model into the side network. In this section, we assess how different ratios of these important tokens affect the overall performance and efficiency. We provide a detailed analysis of FPT+ under varying token ratios, focusing on both 512 × 512 and 256 × 256 input resolutions. We notice that the optimal ratio of important tokens varies across different datasets. For instance, although both DDR and Messidor-2 are fundus image datasets, they show peak performance at different important token ratios. Although a higher ratio of important tokens can provide more information for fusion modules, it carries a risk of misleading the fusion modules and potentially increasing the difficulty of learning. Most datasets prefer a token ratio lower than 0.5, except for Messidor-2 and CHNCXR which achieve the best performance at 0.9 and 0.7, respectively. We conjecture that the optimal token ratio depends on the scale and complexity of the dataset of interest.
In terms of memory efficiency, as the ratio of important tokens increases, the memory consumption also rises dramatically, resulting in significant drops in PME. Across both 512 × 512 and 256 × 256 input resolutions, a token ratio of 0.2 consistently achieves the best trade-off, yielding the highest average AUC while maintaining low memory usage. Although FPT+’s performance may benefit from dataset-specific tuning on important token ratio, a 0.2 token ratio offers a stable and effective choice in a wide range of scenarios.
IV-J Visualization